OpenStudy (samsan9):

Simplify -sqrt(-75) A.) -5i B.) 5i C.) -5isqrt(3) D.) -5isqrt(2)

OpenStudy (imstuck):

\[\sqrt{-75}=5i \sqrt{3}\]and the whole thing is negative, so it is:\[-5i \sqrt{3}\]

OpenStudy (imstuck):

C

OpenStudy (marissalovescats):

@IMStuck you may wanna give a little further explanation on why that is the answer so that they understand and are able to do problems like these on their on in the future.

OpenStudy (samsan9):

i am assuming you just solve it into radical form and add the - at the end :P

OpenStudy (marissalovescats):

Well not really. You have to simplify is first and we know 25*3 is 75 so that simplifies the radical but it's still not negative. So if you remember that i^2=-1 it will give you that negative. \[\sqrt{25*i^2*3} = \sqrt{-75} = -5i \sqrt{3}\] because the square root of i^2 is just i.